problem

A wire 4.00 m long and 6.00 mm in diameter has a r
Answer #1


A)Current= V/R
      =22/0.015
       =1466.67 A
B)
Currrent density=current/area
              =1466.67/(pi*0.003^2)
              =51872722.19 A/m^2
Answer #2

a)

I = V/R = 22/(15X10^-3) = 1.4666 X10^3 A = 1466.66 A


2)

area = 3.14 X(3^2)X(10^-6) m^2 = 28.26 X10^-6


current density = current/area = 1466.66/28.26X10^-6 = 51.898 X 10^6 A/m^2 = 518990327.905 A/m^2